Rogelio’s portfolio is worth 61,000 dollars and has three stocks. It has 18,200 dollars of stock A, which has an expected return of 2.61 percent; it has 1,600 shares of stock B, which has a share price of 8 dollars and an expected return of 8.04 percent; and it has some stock C, which has an expected return of 14.61 percent. What is the expected return of Rogelio’s portfolio? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1
Stock Total value Expected return
A $18200 2.61%
B (1600*8)=$12800 8.04%
C (61000-18200-12800)=$30000 14.61%
Total=$61000

Portfolio return=Respective return*Respective weight

=(18200/61000*2.61)+(12800/61000*8.04)+(30,000/61000*14.61)

=0.0965(Approx)